Three years ago, Juma was three times as old as Ali. In two years time, the sum of their ages will be 62. Determine their present ages
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)


3 years ago
Let Ali's age be x
Juma's age = 3x
2 years later
3+2
x+5+3x+5 =82
4x+10=82
4x=82-10
4x=72
/4
x=8
3 years ago ali was 8 so now he is 11 years
Juma was 3x , which is 24, 24+3=27 years now

Answer by richwmiller(17219)

j-3=3*(a-3)
j+2+a+2=62
j = 58-a
j = 3a-6
58-a = 3a-6
64=4a
a = 16
j = 58-a
j = 42
check
three years ago
13 and 39
two years from now
18+ 44 =62
ok